1. Which one of the following sets of compounds is not characteristic of a buffer? a. HNO3 with NaNO b. NaH POwith Na2HPO: C.please answer 1 and both parts of 2
Answer #1

Solution : 1² Buffer Buffer is. ar mixture of weak auid or weak base and its salt. a. HNO A Naną. HNO is a strong acid and NaBy applying pko= log (18x185) pko = 4.74 Handerson - Hasselbalch equation for buffer. por = pky + tlog [rall] basic [base] роpH= pka + log Isalt] Cacid] pH= 3.74+ log [0:55M] TO EM] Ph: 3:14+ 184 pH: 3.924 ph= 3.92 30, ph=3.92